# Corporate Earnings Calls and Analyst Beliefs

### v1

@@ Abstract
Economic behavior is shaped not only by quantitative information but also by the narratives through which such information is communicated and interpreted (Shiller2017). I show that narratives extracted from earnings calls significantly improve the prediction of both realized earnings and analyst expectations. To uncover the underlying mechanisms, I introduce a novel text-morphing methodology in which large language models generate counterfactual transcripts that systematically vary topical emphasis (the prevailing narrative) while holding quantitative content fixed. This framework allows me to precisely measure how analysts under- and over-react to specific narrative dimensions. The results reveal systematic biases: analysts over-react to sentiment (optimism) and under-react to narratives of risk and uncertainty. Overall, the analysis offers a granular perspective on the mechanisms of expectation formation through the competing narratives embedded in corporate communication.
@@ 1 Introduction
I identify six key narrative dimensions likely to influence analysts’ belief formation: Guidance, Jargon, Confidence, Macro-Perspective, Sentiment, and Uncertainty. These dimensions are selected based on prior research on the effects of linguistic features and stock market reactions. To verify that these narratives are informative, I use a state-of-the-art large language model to assign each earnings call a score from 0 to 10 for the extent to which each narrative dimension is expressed. A ML model trained on these six scores achieves an average out-of-sample performance of approximately relative to models using the full textual embeddings, suggesting that the six narratives capture informational content relevant to analysts’ belief formation.
@@
Most importantly, this in-silico experiment allows me to disentangle the narratives that financial analysts pay attention to—those driving the ML model that predicts their forecasts (Analysts Expectation for Earnings, AEE)—from the narratives they should pay attention to—those associated with future realized earnings (Future Realized Earnings, FRE). I find that, while analysts appear to allocate a rational level of attention to jargon, they severely under-react to narratives of uncertainty (Average PTE for AEE = -9.22 bps versus Average PTE for FRE = -41.09 bps), confidence (Average PTE for AEE = 6.09 bps, Average PTE for FRE = 25.76 bps), and future guidance (Average PTE for AEE = 10.84 bps, Average PTE for FRE = 19.43 bps). Conversely, they over-react to narratives of sentiment (Average PTE for AEE = 34.88 bps, Average PTE for FRE = 25.76 bps) and macro-perspective (Average PTE for AEE = 10.84 bps, Average PTE for FRE = -19.43 bps).
@@
Recent research has begun to use sell-side equity analyst reports as a key data source. DeRosa2024, Ke2024, and Bastianello2024 combine these reports with IBES forecast data to study analysts’ memory and attention, and the structure of their mental models, respectively.
In this paper, I instead examine equity analyst expectation-formation processes using corporate earnings calls, which allows me to measure how an exogenous source of information shapes analysts’ beliefs.
@@
Recent work shows that large language models can reliably reproduce classic behavioral patterns, simulate experimental decision-making environments, and act as transparent proxies for human judgment (Horton2023, Hansen2024, Manning2024, Bhagwat2025, Ludwig2025, Kazinnik2025). They can also implement causal inference through textual interventions, emulate survey-based expectation-formation exercises, and capture heterogeneity in how different “agents” interpret the same information. Together, these findings demonstrate the potential of LLMs to complement or substitute traditional experimental and survey methods in social science research.
@@ 3.1 Analyst Expectation Formation around Earnings Calls
During the earnings call, the newly realized earnings becomes public and enters analysts’ information sets. As a naïve benchmark—abstracting from any private information—there is no reason to expect the future earnings to deviate from the most recently realized level. To characterize departures from this benchmark, 1 defines the Expected Change in Earnings, capturing how analysts believe the future value will differ from the current one. For each firm and horizon , I define
Individual financial analysts naturally form their own expectations about future earnings, even when exposed to the same information during a corporate earnings call. Such dispersion can arise from heterogeneous priors, risk tolerances, models, and, crucially, differences in how the same language is interpreted. Therefore, I examine cross-analyst disagreement using the dispersion of expected changes, defined in Equation 2 for a firm and a forecast horizon , and test whether specific linguistic features are systematically associated with higher or lower disagreement in expectations. To ensure comparability across firms and consistency with other variables here defined, I scale the cross-sectional dispersion of forecasts by the stock price prevailing at the earning call date.
Although corporate earnings calls do not mechanically determine these realizations, I estimate a model predicting the realized change in earnings to test whether the linguistic features of the calls that are systematically associated with realized outcomes differ from those that predict analysts’ behavior, in a later analysis.
@@ 4.1 Predicting Analyst Beliefs with Machine Learning
For each forecast horizon and target variable, the most recent 30% of observations are held out as an out-of-sample test set, while the remaining 70% are used for model fitting and hyper-parameter tuning. To avoid overfitting and to ensure robustness of the selected hyper-parameters, I implement -fold cross-validation within the training set. The empirical cumulative distribution functions of the outcome variables are reported in Figures 10–12 in the Appendix. The training and test sets display closely aligned distributions across horizons.
@@ 4.3 Measuring the Impact of Fundamental News
I compute the interquartile shift, a ”typical” move in SUE, (from the 25th to the 75th percentile) in standardized unexpected earnings (SUE) as the baseline variation. These results provide a benchmark against which, in later sections, I compare the effects of language morphing with large language models, thereby assessing whether earnings call framing can move forecasts in ways comparable to fundamental earnings news.
@@ 5 Counterfactual Earnings Calls with Large Language Models
For each document, I generate the output paragraph by paragraph in sequential forward passes, ensuring internal consistency and preserving global structure. To control the realism and interpretability of the generated texts, I apply prompt engineering techniques that guide the model to preserve numerical content (e.g., financial figures) while varying qualitative language. Specifically, I structure prompts in a way that keeps key values unchanged and induces linguistic transformations within a narrow range of word count variation, thereby enabling controlled comparisons between original and modified statements. See Section C in the appendix for further details.
@@ 5.2 LLM as a Judge
I discard every variant that fails to meet the highest quality standard, thereby ensuring that the predictive analyses are based exclusively on precisely controlled linguistic change. The details on the prompt I use are in Section D.
@@
5.3 Counterfactual Earnings Call Generation
@@
The goal of this exercise is to assess how much of the analyst reaction can be attributed to style and delivery—rather than to substance. More in detail, I generate versions of the same earnings call that vary according to a number of linguistic characteristics, and evaluate how a classifier trained on historical data responds to each variant. The linguistic characteristics I use are:
Confidence: Refers to how assertive and in-control the speaker sounds. This is adjusted by prompting the model to adopt the tone of a confident CEO, using decisive language, avoiding hedging, and conveying authority over strategy and performance. (Mayew2012; DelaParra2024)
Global Focus: Captures the emphasis placed on broader macroeconomic and industry-wide conditions. The model is guided to highlight external forces, such as inflation, regulation, or global trends, and to link them explicitly to the firm’s performance. (Song2025; Link2023; Hassan2019)
Guidance: Reflects how specific and forward-looking the company’s statements are. Stronger guidance is elicited by prompting clear, actionable expectations, directional indicators, and strategic intent, avoiding vague or generic phrasing. (Call2024; Bozanic2018; Anilowski2007)
Sentiment: Denotes the overall emotional tone of the message. The text is modified to sound more positive and optimistic by emphasizing momentum, wins, and enthusiasm for the company’s future. (Price2012; Huang2014)
Jargon: Indicates the degree of technical or finance-specific language used. Higher jargon is introduced through prompts that incorporate institutional and domain-specific terminology without changing the underlying facts.(Li2008; Miller2010)
Uncertainty: Reflects the visibility and emphasis of risk-related language. To increase risk salience, prompts guide the model to explicitly highlight uncertainties, market challenges, and operational risks in a measured and professional tone. (Kravet2013; Lyle2023)
Because the generated calls differ only in their language, any change in the model’s prediction can be interpreted as a quasi-causal effect of linguistic framing on analyst perception. This setup allows me to quantify the persuasive power of narrative, assess the potential for strategic disclosure, and explore whether firms could influence analyst expectations through subtle shifts in communication—without altering the underlying message related to financial fundamentals.
@@
I illustrate this morphing exercise in Figure 7, where I project both the original and morphed transcripts into a two-dimensional semantic space. Each panel corresponds to a distinct linguistic attribute: guidance strength, jargon density, speaker confidence, macro focus, sentiment, and risks mentioned, the original baseline in the bottom panel. The heatmaps reveal that morphing pushes transcripts toward the extremes of these attributes, generating clearly differentiated clusters. This demonstrates that the procedure succeeds in producing systematically varied earnings calls, thereby enabling the in-silico experimental design.
To further illustrate the results of my morphing exercise, I report Table 6, which illustrates how the original text of an earnings call can be systematically morphed along distinct qualitative dimensions. Each row presents a representative passage from an original managerial statement alongside a transformed version that amplifies a specific linguistic or tonal feature. The six dimension i.e., Guidance, Jargon, Confidence, Global Focus, Sentiment, and Uncertainty, capture complementary aspects of managerial communication that may influence analysts’ interpretation of firm fundamentals. These examples serve to clarify the nature of the textual transformations applied in the empirical analysis and to make the corresponding treatments intuitively interpretable.
@@ 5.4 Predicted Treatment Effects (PTEs) induced by Narratives
which measures the change in the model’s predicted outcome when the textual representation is altered to highlight the specified theme, holding all other market-based firm-level and accounting predictors () constant.
@@
Figures 8 and 9 present the average predicted treatment effects (PTEs), expressed in basis points, for each linguistic morph applied to earnings calls. These effects quantify how shifts in communication style—independent of any change in firm fundamentals—alter analysts’ forecasts and realized outcomes. To benchmark the magnitude of these effects, I also report the model-implied response to an interquartile change in fundamental earnings news, defined as the difference in predicted outcomes when standardized unexpected earnings (SUE) move from the 25th to the 75th percentile of its distribution, as computed in prior sections.
@@ 6 Conclusion
This methodology allows me to precisely compute analysts’ over- and under-reaction to a given narrative. I find that financial analysts severely under-react to the narrative of uncertainty and over-react to the optimistic narrative and to macro-perspective framing. I also find that disagreement among analysts is mainly driven by the narrative of uncertainty and macro-perspective framing.
@@ 8 Tables
| (A) Expected Change in Earnings | |||
| Horizon | R-squared (Numeric) | R-squared (Numeric + Text) | Model Improvement (%) |
@@
| (B) Forecast Disagreement | |||
| Horizon | R-squared (Numeric) | R-squared (Numeric + Text) | Model Improvement (%) |
@@
| (C) Realized Change in Earnings | |||
| Horizon | R-squared (Numeric) | R-squared (Numeric + Text) | Model Improvement (%) |
@@
| (A) Expected Change in Earnings | ||
@@
| (C) Realized Change in Earnings | ||
@@
| Morph Type | Original Excerpt | Morphed Excerpt |
@@
| Jargon | We’ve streamlined operations to better align production with customer demand. | We’ve optimized our end-to-end operational footprint through agile capacity realignment to ensure cross-functional synergies across key demand verticals. |
@@
| Uncertainty | We’re encouraged by our progress and expect continued momentum next quarter. | While we’re encouraged by our progress, we remain cautious about potential headwinds–including raw material inflation, geopolitical tensions, and slower–than-expected demand recovery in Europe. |
@@
| Confidence | We think our new product line will perform well, though it’s still early in the rollout. | Our new product line is already exceeding expectations, and we’re absolutely confident it will continue to drive strong performance as adoption scales. |
@@ Appendix A Identification of the latest version of earnings call transcripts
Appendix B Comparison of train/test sample dataset
Appendix C Prompt Descriptions for Language Generation
@@
Appendix D Prompt Descriptions for LLM-as-a-Judge

### v2

@@ Abstract
Economic behavior is shaped not only by quantitative information but also by the narratives through which such information is communicated and interpreted (Shiller2017). I show that narratives extracted from earnings calls significantly improve the prediction of both realized earnings and analyst expectations. To uncover the underlying mechanisms, I introduce a novel text-morphing methodology in which large language models generate counterfactual executive presentation from earnings calls that systematically vary topical emphasis (the prevailing narrative) while holding quantitative content fixed. This framework allows me to precisely measure how analysts under- and over-react to specific narrative dimensions. The results reveal systematic biases: analysts over-react to sentiment (optimism) and under-react to narratives of risk and uncertainty. Overall, the analysis offers a granular perspective on the mechanisms of expectation formation through the competing narratives embedded in corporate communication.
@@ 1 Introduction
I identify six key narrative dimensions likely to influence analysts’ belief formation: Guidance, Jargon, Confidence, Global Perspective, Sentiment, and Uncertainty. These dimensions are selected based on prior research on the effects of linguistic features and stock market reactions. To verify that these narratives are informative, I use a state-of-the-art large language model to assign each earnings call a score from 0 to 10 for the extent to which each narrative dimension is expressed. A ML model trained on these six scores achieves an average out-of-sample performance of approximately relative to models using the full textual embeddings, suggesting that the six narratives capture informational content relevant to analysts’ belief formation.
@@
Most importantly, this in-silico experiment allows me to disentangle the narratives that financial analysts pay attention to—those driving the ML model that predicts their forecasts (Analysts Expectation for Earnings, AEE)—from the narratives they should pay attention to—those associated with future realized earnings (Future Realized Earnings, FRE). I find that, while analysts appear to allocate a rational level of attention to jargon, they severely under-react to narratives of uncertainty (Average PTE for AEE = -9.22 bps versus Average PTE for FRE = -41.09 bps), confidence (Average PTE for AEE = 6.09 bps, Average PTE for FRE = 25.76 bps), and future guidance (Average PTE for AEE = 10.84 bps, Average PTE for FRE = 19.43 bps). Conversely, they over-react to narratives of sentiment (Average PTE for AEE = 34.88 bps, Average PTE for FRE = 25.76 bps) and global perspective (Average PTE for AEE = 10.84 bps, Average PTE for FRE = -19.43 bps).
@@
Recent work has increasingly used sell-side equity analyst reports as a primary data source. For example, DeRosa2024, Ke2024, and Bastianello2024 combine these reports with I/B/E/S forecast data to study different dimensions of analysts’ information processing, including memory, attention, and the structure of their mental models.
In this paper, I take a complementary approach and examine analysts’ expectation-formation processes using corporate earnings calls, which allows me to measure how the information environment of the corporate earnings calls shapes analysts’ beliefs.
@@
Recent work shows that large language models can reliably reproduce classic behavioral patterns, simulate experimental decision-making environments, and act as transparent proxies for human judgment (Horton2023, Hansen2024, Manning2024, Bhagwat2025, Ludwig2025, Kazinnik2025). They can also implement causal inference through textual interventions, emulate survey-based expectation-formation exercises, and capture heterogeneity in how different in-silico “agents” interpret the same information. Together, these findings demonstrate the potential of LLMs to complement or substitute traditional experimental and survey methods in social science research.
@@ 3.1 Analyst Expectation Formation around Earnings Calls
During the earnings call, the newly realized earnings becomes public and enters analysts’ information sets. As a naive benchmark—abstracting from any private information—there is no reason to expect the future earnings to deviate from the most recently realized level. To characterize departures from this benchmark, Eq. 1 defines the Analysts’ Expected Change in Earnings, capturing how analysts believe the future value will differ from the current one. For each firm and horizon , I define
Individual financial analysts naturally form their own expectations about future earnings, even when exposed to the same information during a corporate earnings call. Such dispersion can arise from heterogeneous priors, risk tolerances, models, and, crucially, differences in how the same language is interpreted. Therefore, I examine cross-analyst disagreement using the dispersion of expected changes, defined in Equation 2 for a firm and a forecast horizon , and test whether specific linguistic features are systematically associated with higher or lower disagreement in expectations. To ensure comparability across firms and consistency with other variables here defined, I scale the cross-sectional dispersion of forecasts by the stock price prevailing at the earning call date and define Forecast Disagreement as
Although corporate earnings calls do not mechanically determine these realizations, I estimate a model predicting the realized change in earnings to identify which narratives in corporate communications are most closely related to realized fundamentals. This, in turn, allows me to construct an in-silico counterfactual through which I can study how analysts form their beliefs. I define Realized Change in Earnings as follows:
@@ 4.1 Predicting Analyst Beliefs with Machine Learning
For each forecast horizon and target variable, the most recent 30% of observations are held out as an out-of-sample test set, while the remaining 70% are used for model fitting and hyper-parameter tuning. To avoid overfitting and to ensure robustness of the selected hyper-parameters, I implement -fold cross-validation within the training set.
@@ 4.3 Measuring the Impact of Fundamental News
I compute the inter-quartile shift, a ”typical” move in SUE, (from the 25th to the 75th percentile) in standardized unexpected earnings (SUE) as the baseline variation. These results provide a benchmark against which, in later sections, I compare the effects of language morphing with large language models, thereby assessing whether earnings call framing can move forecasts in ways comparable to fundamental earnings news.
@@ 5 Counterfactual Earnings Calls with Large Language Models
For each document, I generate the output paragraph by paragraph in sequential forward passes, ensuring internal consistency and preserving global structure. To control the realism and interpretability of the generated texts, I apply prompt engineering techniques that guide the model to preserve numerical content (e.g., financial figures) while varying qualitative language. Specifically, I structure prompts in a way that keeps key values unchanged and induces linguistic transformations within a narrow range of word count variation, thereby enabling controlled comparisons between original and modified statements. See Section B in the appendix for further details.
@@ 5.2 LLM as a Judge
I discard every variant that fails to meet the highest quality standard, thereby ensuring that the predictive analyses are based exclusively on precisely controlled linguistic change. The details on the prompt I use are in Section C.
@@
5.3 Narratives and Counterfactual Earnings Calls
@@
The goal of this exercise is to assess how much of the analyst reaction can be attributed to style and delivery, rather than to substance. More in detail, I generate versions of the same earnings call that vary according to a number of linguistic characteristics, and evaluate how a classifier trained on historical data responds to each variant. The linguistic characteristics I use are:
Confidence: Refers to how assertive and in-control the speaker sounds. This is adjusted by prompting the model to adopt the tone of a confident CEO, using decisive language, avoiding hedging, and conveying authority over strategy and performance.
(Mayew2012; DelaParra2024)
Global Focus: Captures the emphasis placed on broader macroeconomic and industry-wide conditions. The model is guided to highlight external forces, such as inflation, regulation, or global trends, and to link them explicitly to the firm’s performance.
(Song2025; Link2023; Hassan2019)
Guidance: Reflects how specific and forward-looking the company’s statements are. Stronger guidance is elicited by prompting clear, actionable expectations, directional indicators, and strategic intent, avoiding vague or generic phrasing.
(Call2024; Bozanic2018; Anilowski2007)
Sentiment: Denotes the overall emotional tone of the message. The text is modified to sound more positive and optimistic by emphasizing momentum, wins, and enthusiasm for the company’s future.
(Price2012; Huang2014)
Jargon: Indicates the degree of technical or finance-specific language used. Higher jargon is introduced through prompts that incorporate institutional and domain-specific terminology without changing the underlying facts.
(Li2008; Miller2010)
Uncertainty: Reflects the visibility and emphasis of risk-related language. To increase risk salience, prompts guide the model to explicitly highlight uncertainties, market challenges, and operational risks in a measured and professional tone.
(Kravet2013; Lyle2023)
Because the generated calls differ only in their language, any change in the model’s prediction can be interpreted as an effect of linguistic framing on analyst perception. This setup allows me to quantify the persuasive power of narrative, assess the potential for strategic disclosure, and explore whether firms could influence analyst expectations through subtle shifts in communication—without altering the underlying message related to financial fundamentals.
@@
I illustrate this morphing exercise in Figure 7, where I project both the original and morphed transcripts into a two-dimensional semantic space. Each panel corresponds to a distinct linguistic attribute: guidance strength, jargon density, speaker confidence, global focus, sentiment, and risks mentioned, the original baseline in the bottom panel. The heatmaps reveal that morphing pushes transcripts toward the extremes of these attributes, generating clearly differentiated clusters. This demonstrates that the procedure succeeds in producing systematically varied earnings calls, thereby enabling the in-silico experimental design.
To further illustrate the results of my morphing exercise, I report Table 6, which illustrates how the original text of an earnings call can be systematically morphed along distinct narrative dimensions. Each row presents a representative passage from an original managerial statement alongside a transformed version that amplifies a specific linguistic or narrative feature. The six dimension i.e., Guidance, Jargon, Confidence, Global Focus, Sentiment, and Uncertainty, capture complementary aspects of managerial communication that may influence analysts’ interpretation of firm fundamentals. These examples serve to clarify the nature of the textual transformations applied in the empirical analysis and to make the corresponding treatments intuitively interpretable.
@@ 5.4 Predicted Treatment Effects (PTEs) induced by Narratives
where is the model trained on the stock characteristics and the original earnings calls- The PTE measures the change in the model’s predicted outcome when the textual representation is altered to highlight the specified theme, holding all other market-based firm-level and accounting predictors () constant.
@@
Figures 8 and 9 present the average predicted treatment effects (PTEs), expressed in basis points, for each linguistic morph applied to earnings calls. These effects quantify how shifts in communication style—independent of any change in firm fundamentals—alter analysts’ forecasts and realized outcomes. To benchmark the magnitude of these effects, I also report the model-implied response to an inter-quartile change in fundamental earnings news, defined as the difference in predicted outcomes when standardized unexpected earnings (SUE) move from the 25th to the 75th percentile of its distribution, as computed in prior sections.
@@ 6 Conclusion
This methodology allows me to precisely compute analysts’ over- and under-reaction to a given narrative. I find that financial analysts severely under-react to the narrative of uncertainty and over-react to the optimistic narrative and to global-focus framing. I also find that disagreement among analysts is mainly driven by the narrative of uncertainty and global-focus perspective.
@@ 8 Tables
| (i) Expected Change in Earnings | |||
| Horizon | R-squared (Fundamentals) | R-squared (Fundamentals + Text) | Gain (%) |
@@
| (ii) Forecast Disagreement | |||
| Horizon | R-squared (Fundamentals) | R-squared (Fundamentals + Text) | Gain (%) |
@@
| (iii) Realized Change in Earnings | |||
| Horizon | R-squared (Fundamentals) | R-squared (Fundamentals + Text) | Gain (%) |
@@
| (i) Expected Change in Earnings | ||
@@
| (iii) Realized Change in Earnings | ||
@@
| Narrative | Original Excerpt | Morphed Excerpt |
@@
| Jargon | We have streamlined operations to better align production with customer demand. | We’ve optimized our end-to-end operational footprint through agile capacity realignment to ensure cross-functional synergies across key demand verticals. |
@@
| Uncertainty | We are encouraged by our progress and expect continued momentum next quarter. | While we are encouraged by our progress, we remain cautious about potential headwinds–including raw material inflation, geopolitical tensions, and slower–than-expected demand recovery in Europe. |
@@
| Confidence | We think our new product line will perform well, though it is still early in the rollout. | Our new product line is already exceeding expectations, and we are absolutely confident it will continue to drive strong performance as adoption scales. |
@@ Appendix A Identification of the latest version of earnings call transcripts
Appendix B Prompt Descriptions for Language Generation
@@
Appendix C Prompt Descriptions for LLM-as-a-Judge